OpenStudy (anonymous):

In a circle whose radius is 4 centimeters, what is the length, in centimeters, of an arc intercepted by a central angle of 2 1/2 radians?

OpenStudy (mathstudent55):

The circumference of a cirdcle is \( C = 2 \pi r \) This is based on the entire circle being an angle of \(2 \pi\) radians. For the length of an arc of a circle with a central angle of \( \theta\), the angle is the following fraction of the the circle \(\dfrac{\theta}{2 \pi} \). The length of an arc with cetral angle of \( \theta\) is \(s = \dfrac{\theta}{2 \pi} \times 2 \pi r = r \theta \) All you need to do is use the formula: \(s = r \theta \) where s = arc length, r = radius, \(\theta\) = central angle.